2. a) Find the derivatives of the following functions with respect to x.


X^3+Y^3=3



Y= (sin x) ^tan x

1
Expert's answer
2021-12-09T18:47:18-0500

Q1)

x³ + y³ = 3

The function is in implicit form.

Differentiating with respect to x

3x² + 3y² "\frac{dy}{dx}=0"

=> 3y² "\frac{dy}{dx}" = -3x²

=> y² "\frac{dy}{dx}=" -x²

=> "\frac{dy}{dx}= -\frac{x²}{y²}"

Q2

y = (sin x)tan x

Taking logarithm of both sides

logey = loge(sin x)tan x

=> logey =tan x loge(sin x)

Differentiating with respect to x

"\frac{1}{y}\frac{dy}{dx}= tanx\frac{d}{dx}log_{e}sinx+log_{e}sinx\frac{d}{dx}tanx"

"\frac{1}{y}\frac{dy}{dx}= tanx\frac{1}{sinx}cosx +log_{e}sinx.(sec²x)"

=> "\frac{1}{y}\frac{dy}{dx}= tanx.cotx +log_{e}sinx.(sec²x)"

=> "\frac{dy}{dx}" = y (1 +sec²x logesinx)

=> "\frac{dy}{dx}" = (sin x)tan x(1 + sec²x logesinx)